# MARM: Unlocking the Future of Recommendation Systems through Memory Augmentation and Scalable Complexity

### v2

@@ 4. Conclusion
We propose the Memory Augmented Recommendation Model (MARM), a multi-layer recommendation model that utilizes Memory Cache to accelerate inference. In recommendation scenarios, computational complexity is a significant performance constraint. We use caching to store partial computation results of complex models, reducing the complexity of single modeling from $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$. Based on the MARM framework, we can extend sequence modeling from single-layer target attention to multi-layer with linear resource consumption, significantly breaking through the computational bottleneck of recommendation models and enabling modeling of users’ life-long histories. We explored the scaling law within the MARM framework, confirming a proportional relationship between cache size and recommendation performance. Furthermore, our MARM method can seamlessly integrate with existing recommendation models, including fine ranking, coarse ranking, and retrieval. Through the MARM series of methods, we achieved 2.079% increase in average app viewing time per user in Kuaishou’s overseas operations, demonstrating the effectiveness and practicality of the MARM model. To the best of our knowledge, MARM is the first recommendation model to achieve multi-dimensional scaling in the field of life-long sequence modeling.

### v3

@@ 5. Conclusion
We propose the MARM, a multi-layer recommendation model that utilizes Memory Cache to accelerate inference. In recommendation scenarios, computational complexity is a significant performance constraint. We use caching to store partial computation results of complex models, reducing the complexity of single modeling from $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$. Based on the MARM framework, we can extend sequence modeling from single-layer target attention to multi-layer with linear resource consumption, significantly breaking through the computational bottleneck of recommendation models and enabling modeling of users’ life-long histories. We explored the scaling laws within the MARM framework, confirming a proportional relationship between cache size and recommendation performance. Furthermore, our MARM method can seamlessly integrate with existing recommendation models, including fine ranking, coarse ranking, and retrieval.
6. GenAI Usage Disclosure
In this paper, we only utilize the AI tools to fix our grammar mistakes. The research motivation, methodology, and experimental results are entirely derived from first-hand experiments and analysis conducted in real-world business scenarios. All data and observations have been rigorously validated by online A/B tests and offline analyses.